OpenStudy (anonymous):

Y=mx+b define which variable is the y intercept and define which one is the x intercept

OpenStudy (anonymous):

b is the y intercept

OpenStudy (anonymous):

m is the slope

OpenStudy (anonymous):

0 = mx + b -b = mx x = -b/m x intercept is -b/m

OpenStudy (anonymous):

so we have to find the slope intercept of 8x-10y=-100 what is the y intercept and wht is the x intercept

OpenStudy (anonymous):

move the numbers around so it's -10y=8x-100. then divide the right side by -10. simplify it so it's y= (-4/5)x+10

OpenStudy (anonymous):

from there use what moneybird showed you by comparing it to the y=mx+b formula.

OpenStudy (anonymous):

so the y intercept is 10 then what is the x intercept? there is no number so will it be 1?

OpenStudy (anonymous):

no it's 10 divided by -4/5

OpenStudy (anonymous):

x intercept is either -b/m or put y = 0 and figure x out that way.

OpenStudy (anonymous):

and change it to a negative

OpenStudy (anonymous):

its negative 10 divided by negative 4/5 so it ends up positive.

OpenStudy (anonymous):

damn u moneybird lol